Kumar, Inc., evaluates managers of producing departments on their ability to control costs. In addition to the costs directly traceable to their departments, each production manager is held responsible for a share of the costs of a support center, the Human Resources (HR) Department. The total costs of HR are allocated on the basis of actual direct labor hours used. The total costs of HR and the actual direct labor hours worked by each producing department are as follows:

Year 1 Year 2
Direct labor hours worked:
Department A 24,000 27,000
Department B 36,000 27,000
Total hours 60,000 54,000
Actual HR cost $160,000 $160,000
Budgeted HR cost 115,000 * 113,500 *

*$0.25 per direct labor hour plus $100,000.

Required:

Allocate the HR costs to each producing department for Year 1 and Year 2 using the direct method with actual direct labor hours and actual HR costs.

Year 1 Year 2
Department A $ $
Department B $ $
Answer #1

Year 1:

Department A = $160,000 X 24,000 / 60,000 = $64,000

Department B = $160,000 X 36,000 / 60,000 = $96,000

Year 2:

Department A = $160,000 X 27,000 / 54,000 = $80,000

Department B = $160,000 X 27,000 / 54,000 = $80,000

Year 1 Year 2
Department A $64,000 $80,000
Department B $96,000 $80,000